OpenStudy (anonymous):

Helpp Simplyfing Rational Exponets 3x^-1/2*3x^1/2y^-1/3 /3y^-7/4 (m*m^-2 n^5/3)^2

OpenStudy (anonymous):

does anyone get it??

OpenStudy (saifoo.khan):

Please rewrite. :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[3x^{-1/2}\times3x ^{1/2}y ^{-1/3}\div3y ^{-7/4}\]

OpenStudy (anonymous):

\[(m*m ^{-2}n ^{5/3})^{2}\]

OpenStudy (anonymous):

i rewrote it :)))

OpenStudy (saifoo.khan):

Which one is the question?

OpenStudy (anonymous):

both

OpenStudy (saifoo.khan):

\[\Huge 3y^\frac{17}{12}\]

OpenStudy (anonymous):

how did u get that

OpenStudy (saifoo.khan):

For the first one., \[\Huge 3 \times 3 \times \frac{1}{3}* x^{-\frac12+\frac12} * y^{-\frac13 + \frac74}\]

OpenStudy (anonymous):

its negative 7/4 not positive

OpenStudy (saifoo.khan):

\[\Huge \div 3y^{-\frac17} \to \times \frac{1}{3y^{-\frac17}} \to 3y^\frac17\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

okay i get it but, the next problem i had a little trouble with the m's

OpenStudy (saifoo.khan):

\[\Large (m*m ^{-2}n ^{5/3})^{2} \to (m^{-1}n^\frac53)^2 \to (m^{-2}n^\frac{10}{3}) \to (\frac{n^\frac{10}{3}}{m^2})\]